### v3

@@ Abstract
Decision-making is an essential attribute of any intelligent agent or group. Natural systems are known to converge to optimal strategies through at least two distinct mechanisms: collective decision-making via imitation of others, and individual trial-and-error. This paper establishes an equivalence between these two paradigms by drawing from the well-established collective decision-making model of nest-site selection in swarms of honey bees. We show that the emergent distributed cognition (sometimes referred to as the hive mind) arising from individual bees following simple, local imitation-based rules is equivalent to a single online reinforcement learning (RL) agent interacting with many parallel environments. The update rule through which this macro-agent learns is a bandit algorithm that we coin Maynard-Cross Learning. Our analysis implies that a group of cognition-limited organisms can be on-par with a more complex, reinforcement-enabled entity, substantiating the idea that group-level intelligence may explain how seemingly simple and blind individual behaviors are selected in nature.
@@ 1 Introduction
of $N$-bees
[…]
@@
Figure 1: The “hive mind” of a swarm of $N$ bees selecting a nest site among $n$ options is a single $n$-armed bandit RL agent learning from $N$ environments in parallel.
@@
Decision-making is the ability to choose the best action amongst all available options in a given scenario. Defining what a “best action” means requires comparing all possible outcomes [VonNeumann]. Most often, the preference among outcomes is abstracted out as a single scalar signal [silver2022] conveyed to the agent/individual [^1] To avoid confusion, we use “agent” in the context of RL and “individual” in the context of a population. by the environment as feedback for taking a decision, which casts optimal decision-making into a maximization problem. In nature, at least two distinct mechanisms allow single agents or groups of individuals to converge towards making optimal decisions.
@@
The first of these mechanisms adopts an individualistic approach, whereby the agent learns to take optimal decisions through trial and error. This paradigm is known as Reinforcement Learning (RL), where the agent learns a policy that maximizes the rewards it receives upon interaction with the environment [sutton2018reinforcement]. Numerous studies in computational neuroscience have established links between learning processes happening in the brains of living beings and the formal framework of algorithmic RL [neftci2019reinforcement, Muller2024]. In this paper, we are specifically interested in the multi-armed bandit framework [sutton2018reinforcement], where a single RL agent learns to make choices among different options (or “arms”) in an online fashion. Online RL agents learn by either interacting with a single environment (a setting recently referred to as the “streaming” setting [elsayed2024streamingdeepreinforcementlearning, vasan2024deeppolicygradientmethods]), or, when possible, by collecting samples in parallel from multiple copies of the same environment simultaneously (“parallel” setting). To stabilize training, the learning rate and the number of parallel environments are crucial in the streaming and parallel cases, respectively. Among the many learning algorithms designed to solve multi-armed bandits (Upper-Confidence-Bound [auerUCB1], $\epsilon$-greedy [sutton2018reinforcement], Gradient Bandit [Williams1992], etc.), we consider the Cross Learning (CL) [cross1973] update rule (named after the economist John G. Cross [bloembergen2015evolutionary] and not to be mixed with more recent overlapping nomenclature). CL is closely related to the Gradient Bandit algorithm, an RL algorithm directly based on the policy gradient theorem [sutton2018reinforcement]. The policy gradient theorem is also the basis of many more advanced policy gradient (e.g., REINFORCE [Williams92] and PPO [schulman2017]) and actor-critic (e.g., SAC [haarnoja2018]) RL algorithms.
@@
The second mechanism takes a collective approach, where individuals mimic other individuals in the group to make decisions. Typically, this process is cast as a consensus problem, in which all individuals in a population have to agree on the best decision from a set of alternatives. This problem is commonly referred to as the best-of-n collective decision-making problem (CDM). CDM arises in various domains, such as honey bee colonies [Reina15, bose2017], human societies [Jackson10], and robot swarms [valentini2016collective, valentini2014self]. Imitation of success [sandholm2010population], from Evolutionary Game Theory (EGT), is a solution to CDM where individuals mimic randomly chosen neighbors based on how successful those neighbors’ decisions have been. Another solution to CDM is the weighted voter model [valentini2014self, Reina_2024], which comes from modeling the house-hunting behavior of honey bees [scott_99]. During nest-site selection, honey bees are known to adopt the behavior represented in Fig. 1 (left-hand half): after scouting one of $n$ potential nesting areas, bees come back to the swarm and perform a “waggle dance” [scott_99] that describes the coordinates of the option they have explored. This dance is performed at a frequency that is proportional to the estimated quality of the explored area. Other bees go scout the area corresponding to the first dance they witness, and this process repeats until the colony reaches a quorum [seeley2004quorum]. At that point, the entire swarm takes off and leaves for the winning site (this is a simplified description of bees’ behavior, see [seeley2004group, seeley2004quorum, passino2006modeling, Seeley08] for details). We refer to this distributed behavior as that of a “hive mind”, motivated by two factors: (1) biologists have already witnessed parallels between the emergent distributed cognition arising from individual-to-individual interactions and cognition in neuron-based complex brains of vertebrates [Seeley08], and (2): we formally show that the bee colony collectively learns as if it were one single RL entity.
@@
In this paper, we highlight how imitation mechanisms in groups and individual trial-and-error mechanisms can be bridged via different variants of the Replicator Dynamic [sandholm08] (RD). In particular, we show that the model used by biologists to describe the house-hunting behavior of honey bees amounts to a single-agent RL process at the swarm level, as illustrated in Fig. 1. To the best of our knowledge, our work is the first to formally establish this equivalence. Specifically, our contributions are:
@@ 2 Background
Multi-armed bandits (Fig. 1, right-hand half) are the simplest type of environments encountered in RL literature. They consist of a discrete set of available actions, called “arms”, amongst which the agent has to find the most rewarding. In the $n$-armed bandits considered in this paper, pulling an arm $a\in\{1,\dots,n\}$ returns a real-valued reward $r_{a}\in[0,1]$ sampled from a hidden distribution $r(a)$. The objective for an RL agent playing a multi-armed bandit is to learn a policy, denoted by the probability vector $\pi=(\pi_{1},\dots.\pi_{n})$, that maximizes the rewards obtained upon pulling the arms. Different optimization strategies exist to find such policies, one of the oldest being Cross Learning [cross1973]:
Definition 2.1.
@@ 2.2 Evolutionary Game Theory
Evolutionary Game Theory (EGT) is the study of population games [sandholm2010population]. A population $\mathcal{P}$ is made of a large number of individuals, where any individual $i$ is associated with a type, denoted by $T_{i}\in\{1,\dots,n\}$. The population vector $\pi=(\pi_{1},\dots,\pi_{n})$ represents the fraction of individuals in each type ($\sum_{i}\pi_{i}=1$). Individuals are repeatedly paired at random to play a game, each receiving a separate payoff. Individuals adapt their type based on their payoff according to a revision protocol.
Remark 2.2.
Population-policy equivalence. For the argument of our paper, it is interesting to interchangeably define $\pi=(\pi_{1},\dots,\pi_{n})$ both as a multi-armed bandit RL policy and as a population vector [bloembergen2015evolutionary]. This is possible because in both cases the vector $\pi$ is constrained to the probability simplex. Further, note that uniformly sampling an individual of type $a$ from the population $\mathcal{P}$ (represented by the population vector $\pi$) is equivalent to sampling an action $a$ from the policy $\pi$.
2.2.1 Imitation of success and TRD
In evolutionary dynamics, an important revision protocol is imitation of success [sandholm2010population]:
Definition 2.3.
@@ 2.2.1 Imitation of success and TRD
One can easily see why this rule is called “imitation of success”: $i$ imitates $j$ based on $j$’s payoff. Imitation of success is typically used to model replication of the fittest in evolution for biology. Moreover, this is also used to model decision-making in economics [apesteguia2007imitation] (for instance to model how consumers trust online reviews to make decisions, etc.). When aggregated across the population, this revision model yields a famous evolutionary dynamic known as the Taylor Replicator Dynamic [taylor78, sandholm08] (TRD) (see Section 3.1):
@@
2.2.2 Weighted Voter Model and MRD
The weighted voter model [Reina_2024] is typically used to model the house-hunting behavior of honey bees [valentini2016collective, scott_99]:
Definition 2.4.
In the “weighted voter” revision model $R_{\text{wvoter}}$, any bee $i\in\mathcal{P}$ of type $T_{i}=a$ (where $a$ corresponds to a nest-site option) executes the following process: $i$ estimates the quality of its current type $r_{a}\sim r(a)$, where $0\leq r_{a}\leq 1$. After obtaining $r_{a}$, $i$ locally broadcasts its type at a frequency proportional to $r_{a}$. $i$ switches its type to the first type $b$ that it perceives from its neighborhood. Assuming all individuals are well mixed in the population [martin06], the corresponding expected probability of $i$ switching to type $b$ is the proportion of votes cast for $b$: $P(b\leftarrow a)=\frac{N_{b}\mathbb{E}[r_{b}]}{\sum_{l}N_{l}\mathbb{E}[r_{l}]}$ where $N_{k}$ is the number of individuals of type $k$. By dividing the numerator and the denominator by the size of the population, this probability can be written : $P(b\leftarrow a)=\frac{\pi_{b}\mathbb{E}[r_{b}]}{\sum_{l}\pi_{l}\mathbb{E}[r_{l}]}$
Note that, in this model, honey bees do not need to directly observe the quality estimate of other scout bees, but only their type. In Section 3.2, we show that the weighted voter revision model aggregates to a variant of the TRD, called the Maynard-Smith Replicator Dynamic [smith82] (MRD):
@@ 3 Methodology
In Section 3.1, we remark that previous results from the literature yield an interesting insight binding $R_{\text{success}}$ and reinforcement learning together: a large population following the “imitation of success” revision model can equivalently be considered as a single RL agent. While this is a direct consequence of previously known results, we could not find this insight clearly formulated in the literature and thus we took the liberty to formalize it as 1. Then, in Section 3.2, we prove that the weighted voter revision model also aggregates to an RL algorithm. In other words, our analysis shows that, at least in the well-studied case of nest-site selection, a swarm of honey bees collectively acts as a single RL entity. We formalize this result as 2.
@@ 3.1 Imitation of Success and Cross Learning
Evolutionary Game Theorists have long been interested in the “imitation of success” revision protocol, as it models replication of the fittest in the evolutionary setting. In this literature, it is famously known that a population of individuals following $R_{\text{success}}$ aggregates to the Taylor Replicator Dynamic (see for instance [sandholm2010population, sandholm08]). A proof of this result can be found in the Appendix: {lemmarep} An infinite population of individuals adopting $R_{\text{success}}$ follows the TRD:
@@
where $\pi_{a}$ is the proportion of type $a$ in the population, $q^{\pi}_{a}$ is the expected payoff (also called “fitness”) of type $a$ against the population, and $v^{\pi}$ is the average fitness of the population. {appendixproof} Let $P(a\leftarrow b)$ denote the inflow of individuals of type $b$ into type $a$, i.e, the proportion of the population leaving type $b$ and adopting type $a$. The population has a proportion of $\pi_{b}$ individuals of type $b$, each having a probability $\pi_{a}$ of meeting an individual of type $a$, and a conditional probability $\mathbb{E}[r_{a}]$ of switching to its type. Thus, we get $P(a\leftarrow b)=\pi_{b}\pi_{a}\mathbb{E}[r_{a}]$:
$d\pi_{a}$ $=\sum_{b\neq a}\underbrace{P(a\leftarrow b)}_{\text{inflow}}-\underbrace{P(b\leftarrow a)}_{\text{outflow}}$ (6)
$=\sum_{b\neq a}\pi_{b}\pi_{a}\mathbb{E}[r_{a}]-\pi_{a}\pi_{b}\mathbb{E}[r_{b}]$
$=\pi_{a}\Bigr{[}\sum_{b\neq a}\pi_{b}\mathbb{E}[r_{a}]-\sum_{b\neq a}\pi_{b}\mathbb{E}[r_{b}]\Bigr{]}\hskip 20.00003pt\sum_{b\neq a}\pi_{b}+\pi_{a}=1$
$=\pi_{a}\Bigr{[}(1-\pi_{a})\mathbb{E}[r_{a}]-\sum_{b\neq a}\pi_{b}\mathbb{E}[r_{b}]\Bigr{]}$
$=\pi_{a}\Bigr{[}\mathbb{E}[r_{a}]-\sum_{b}\pi_{b}\mathbb{E}[r_{b}]\Bigr{]}$
$=\pi_{a}(q^{\pi}_{a}-v^{\pi})$ (7)
Our choice of notation in Section 3.1 is reminiscent of the RL literature and may seem unusual from an EGT perspective. In fact, this choice is motivated by another, perhaps more surprising result from [borgers1997learning, bloembergen2015evolutionary], who showed that the Cross Learning RL algorithm performs updates that are also similar to the TRD (proof in Appendix):
{lemmarep}
In expectation, an RL agent learning via the CL update rule follows:
@@
where $q^{\pi}_{a}$ is the action-value of $a$, and $v^{\pi}$ is the value of policy $\pi$. {appendixproof} Let us compute the expectation over actions sampled from $\pi$ in Eq. 2. For convenience, we write
$\mathbb{E}[d\pi_{a}]:=\mathbb{E}_{k\sim\pi}[d\pi_{a}(k)]$, and $\mathbb{E}[r_{k}]:=\mathbb{E}_{r_{k}\sim r(k)}[r_{k}]$:
$\mathbb{E}[d\pi_{a}]$ $=\sum_{k=1}^{n}\pi_{k}.d\pi_{a}(k)$ (9)
$=\pi_{a}.d\pi_{a}(a)+\sum_{k\neq a}\pi_{k}.d\pi_{a}(k)$
$=\pi_{a}\mathbb{E}[r_{a}](1-\pi_{a})+\sum_{k\neq a}\pi_{k}\mathbb{E}[r_{k}](-\pi_{a})$
$=\pi_{a}\Bigr{[}\mathbb{E}[r_{a}]-\pi_{a}\mathbb{E}[r_{a}]-\sum_{k\neq a}\pi_{k}\mathbb{E}[r_{k}]\Bigr{]}$
$=\pi_{a}\Bigr{[}\mathbb{E}[r_{a}]-\sum_{k}\pi_{k}\mathbb{E}[r_{k}]\Bigr{]}$
$=\pi_{a}(q^{\pi}_{a}-v^{\pi})$ (10)
With this notation, it is straightforward to combine Sections 3.1 and 3.1, provided a sensible duality exists between the identical terms. Remember how Remark 2.2 describes a duality between the population vector of Section 3.1 and the policy of Section 3.1, $\pi$. In Section 3.1, the fitness $q^{\pi}_{a}$ is the payoff that an individual of type $a$ can expect on average when encountering a random individual from the population $\pi$. In Section 3.1, the dual of this individual of type $a$ is an action $a$ sampled from the policy $\pi$, whose expected reward is the action-value $q^{\pi}_{a}$. Similarly, the dual of the population fitness $v^{\pi}$ is the policy-value. In other words, under $R_{\text{success}}$, individuals sampled from the population $\pi$ can equivalently be regarded as action samples from an RL macro-agent, whose rewards are the individuals’ payoffs. The agent maximizes the average payoff of the group via Exact Cross Learning (we call “exact” the RL algorithm that directly applies expected updates instead of updates computed from sample estimates):
@@
Proof 3.1.
Direct consequence of Sections 3.1 and 3.1.
1 intuitively describes situations where individuals have individualistic incentives. But let us now turn our attention to simple, cognition-limited organisms that may not even have a notion of individual success to optimize: honey bees.

### v4

@@ Abstract
Decision-making is an essential attribute of any intelligent agent or group. Natural systems are known to converge to optimal strategies through at least two distinct mechanisms: collective decision-making via imitation of others, and individual trial-and-error. This paper establishes an equivalence between these two paradigms by drawing from the well-established collective decision-making model of nest-hunting in swarms of honey bees. We show that the emergent distributed cognition (sometimes referred to as the hive mind) arising from individual bees following simple, local imitation-based rules is that of a single online reinforcement learning (RL) agent interacting with many parallel environments. The update rule through which this macro-agent learns is a bandit algorithm that we coin Maynard-Cross Learning. Our analysis implies that a group of cognition-limited organisms can be equivalent to a more complex, reinforcement-enabled entity, substantiating the idea that group-level intelligence may explain how seemingly simple and blind individual behaviors are selected in nature.
From a biological perspective, this analysis suggests how such imitation strategies evolved: they constitute a scalable form of reinforcement learning at the group level, aligning with theories of kin and group selection. Beyond biology, the framework offers new tools for analyzing economic and social systems where individuals imitate successful strategies, effectively participating in a collective learning process. In swarm intelligence, our findings will inform the design of scalable collective systems in artificial domains, enabling RL-inspired mechanisms for coordination and adaptability at scale.
@@ 1 Introduction
of NN-bees
$a_{B}\sim$$\mathcal{P}$
$a_{C}\sim$$\mathcal{P}$
$a_{A}\sim$$\mathcal{P}$
$a_{B}\sim$$\pi$
$a_{D}\sim$$\mathcal{P}$
@@
Figure 1: The “hive mind” of a swarm of $N$ bees nest-hunting among $n$ options is a single $n$-armed bandit RL agent learning from NN environments in parallel.
@@
Decision-making is the ability to choose the best action amongst all available options in a given scenario. Defining what a “best action” means requires ranking all possible outcomes [1]. Most often, the preference among outcomes is abstracted out as a single scalar signal [2] conveyed to the agent/individual [^1] To avoid confusion, we use “agent” in the context of RL and “individual” in the context of a population. by the environment as feedback for making a decision, which casts optimal decision-making into a maximization problem.
@@
In nature, at least two distinct mechanisms allow single agents or groups of individuals to converge towards making optimal decisions. The first of these mechanisms adopts an individualistic approach, whereby the agent learns to make optimal decisions through trial and error. This paradigm is known as Reinforcement Learning (RL), where the agent learns a policy that maximizes the rewards it receives upon interacting with its environment [3]. Numerous studies in computational neuroscience have established links between learning processes happening in the brains of living beings and the formal framework of algorithmic RL [4, 5]. In this paper, we are specifically interested in the multi-armed bandit framework [3], where a single RL agent learns to make choices among different options (or “arms”) in an online fashion. Online RL agents learn by either interacting with a single environment (a setting recently referred to as the “streaming” setting [6, 7]), or, when possible, by collecting samples in parallel from multiple copies of the same environment simultaneously (the “parallel” setting [8]). To stabilize RL, the learning rate and the number of parallel environments are crucial in the streaming and parallel cases, respectively. Among the many learning algorithms designed to solve multi-armed bandits (Upper-Confidence-Bound [9], $\epsilon$-greedy [3], Gradient Bandit [10], etc.), we consider the Cross Learning (CL) [11] update rule (named after the economist John G. Cross [12] and not to be mixed with more recent overlapping nomenclature). CL is closely related to the Gradient Bandit algorithm, an RL algorithm directly based on the policy gradient theorem [3]. The policy gradient theorem is also the basis of many more advanced policy gradient (e.g., REINFORCE [13] and PPO [14]) and actor-critic (e.g., SAC [15]) RL algorithms.
@@
The second mechanism takes a collective approach, where individuals mimic other individuals in the group to make decisions. Typically, this process is cast as a consensus problem, in which all individuals in a population have to agree on the best decision from a set of alternatives. This problem is commonly referred to as best-of-n collective decision-making (CDM). CDM arises in various domains, such as honey bee colonies [16, 17], human societies [18], and robot swarms [19, 20]. Imitation of success [21], from Evolutionary Game Theory (EGT), is a solution to CDM where individuals mimic randomly chosen neighbors based on how successful those neighbors’ decisions have been. Another solution to CDM is the weighted voter model [20, 22], which comes from modeling the nest-hunting behavior of honey bees [23]. During nest-hunting, honey bees are known to adopt the behavior represented in Fig. 1 (left-hand half): after scouting one of $n$ potential nesting areas, bees come back to the initial location of the swarm and perform a “waggle dance” [23] that describes the coordinates of the option they have explored. This dance is performed at a frequency that is proportional to the estimated quality of the explored area. Other bees go scout the area corresponding to the first dance they witness, and this process repeats until the colony reaches a quorum [24]. At this point, the entire swarm takes off and leaves for the winning site (note this is a simplified description of bees’ behavior, see [25, 24, 26, 27] for details). We refer to this distributed behavior as that of a “hive mind”, motivated by two factors: (1) biologists have already witnessed parallels between the emergent distributed cognition arising from individual-to-individual interactions and cognition in neuron-based complex brains of vertebrates [27], and (2): we formally show that the bee colony collectively learns as if it were one single RL entity.
@@
In this paper, we highlight how imitation mechanisms in groups and individual trial-and-error mechanisms can be bridged via different variants of the Replicator Dynamic [28] (RD). In particular, we show that the model used by biologists to describe the nest-hunting behavior of honey bees amounts to a single-agent RL process at the swarm level, as illustrated in Fig. 1. To the best of our knowledge, this is the first work to formally establish this equivalence. Specifically, our contributions are:
@@
The first two contributions, which detail how social imitation models aggregate into a single reinforcement learning agent, provide a formal description of diverse instances of collective intelligence.
@@ 2 Background
Multi-armed bandits (Fig. 1, right-hand half) are the simplest type of environment encountered in RL literature. They consist of a discrete set of available actions, called “arms”, among which the agent has to find the most rewarding. In the $n$-armed bandits considered in this paper, pulling an arm $a\in\{1,\dots,n\}$ returns a real-valued reward $r_{a}\in[0,1]$ sampled from a hidden distribution $r(a)$. The objective for an RL agent playing a multi-armed bandit is to learn a policy, denoted by the probability vector $\pi=(\pi_{1},\dots.\pi_{n})$, that maximizes the rewards obtained upon pulling the arms. Different optimization strategies exist to find such policies, one of the oldest being Cross Learning [11]:
Definition 1.
@@ 2.2 Evolutionary Game Theory
Evolutionary Game Theory (EGT) is the study of population games [21]. A population $\mathcal{P}$ is made of a large number of individuals, where any individual $i$ is associated with a type, denoted by $T_{i}\in\{1,\dots,n\}$. The population vector $\pi=(\pi_{1},\dots,\pi_{n})$ represents the fraction of individuals in each type ($\sum_{i}\pi_{i}=1$). Individuals are repeatedly paired at random to play a game, each receiving a separate payoff. Individuals adapt their type based on their payoff according to a revision protocol.
Remark 1.
Population-policy equivalence. For the argument of our paper, it is interesting to interchangeably define $\pi=(\pi_{1},\dots,\pi_{n})$ both as a multi-armed bandit RL policy and as a population vector [12]. This is possible because in both cases the vector $\pi$ is constrained to the probability simplex. Further, note that uniformly sampling an individual of type $a$ from the population $\mathcal{P}$ (represented by the population vector $\pi$) is equivalent to sampling an action $a$ from the policy $\pi$.
2.2.1 Imitation of success and the Taylor Replicator Dynamic
In evolutionary dynamics, an important revision protocol is imitation of success [21]:
Definition 2.
@@ 2.2.1 Imitation of success and the Taylor Replicator Dynamic
One can easily see why this rule is called “imitation of success”: $i$ imitates $j$ based on $j$’s payoff. Imitation of success is typically used to model replication of the fittest in evolution for biology, or certain human behaviors in economics [29]. When aggregated across the population, this revision model yields a famous evolutionary dynamic known as the Taylor Replicator Dynamic [30, 28] (TRD) (see LABEL:lem:PG):
@@
2.2.2 Weighted Voter Model and the Maynard-Smith Replicator Dynamic
The weighted voter model [22] instead typically models the nest-hunting behavior of honey bees described in Section 1 [19, 23]:
Definition 3.
In the “weighted voter” revision model $R_{\text{wvoter}}$, any bee $i\in\mathcal{P}$ of type $T_{i}=a$ (where $a$ corresponds to a nest-site option) executes the following process: $i$ estimates the quality of its current type $r_{a}\sim r(a)$, where $0\leq r_{a}\leq 1$. After obtaining $r_{a}$, $i$ locally broadcasts its type at a frequency proportional to $r_{a}$. $i$ switches its type to the first type $b$ that it perceives from its neighborhood. Assuming all individuals are well mixed in the population [31], the corresponding expected probability of $i$ switching to type $b$ is the proportion of votes cast for $b$ within its neighborhood: $P^{(i)}_{\text{neighborhood}}(b\leftarrow a)=\frac{N^{(i)}_{b}\mathbb{E}[r_{b}]}{\sum_{l}N^{(i)}_{l}\mathbb{E}[r_{l}]}$ where $N^{(i)}_{k}$ is the number of individuals of type $k$ in the neighborhood of $i$.
Note that, in this model, honey bees do not need to directly observe the quality estimate of other scout bees, but only their type. In Section 3.2, we show that the weighted voter revision model aggregates to a variant of the TRD, called the Maynard-Smith Replicator Dynamic [32] (MRD):
@@ 3 Methodology
In Section 3.1, we remark that previous results from the literature yield an interesting insight binding $R_{\text{success}}$ and reinforcement learning together: a large population following the “imitation of success” revision model can equivalently be considered as a single RL agent. While this is a direct consequence of previously known results, we could not find this insight formulated in the literature and thus we took the liberty to formalize it as 1. Then, in Section 3.2, we prove that the weighted voter revision model also aggregates to an RL algorithm. In other words, our analysis indicates that, at least in the well-studied case of nest-hunting, a swarm of honeybees collectively acts as a single RL entity. We formalize this result as 2.
@@ 3.1 Imitation of Success and Cross Learning
Evolutionary Game Theorists have long been interested in the “imitation of success” revision protocol, as it models replication of the fittest in the evolutionary setting. In this literature, it is famously known that a population of individuals following $R_{\text{success}}$ aggregates to the Taylor Replicator Dynamic (see for instance [21, 28] and the proof in the Appendix):
Lemma 1.
lem:PG An infinite population of individuals adopting $R_{\text{success}}$ follows the TRD:
@@
where $\pi_{a}$ is the proportion of type $a$ in the population, $q^{\pi}_{a}$ is the expected payoff (also called “fitness”) of type $a$ against the population, and $v^{\pi}$ is the average fitness of the population.
Our choice of notation in LABEL:lem:PG is reminiscent of the RL literature and may seem unusual from an EGT perspective. In fact, this choice is motivated by another, lesser-known result from [33, 12], who showed that the Cross Learning RL algorithm performs updates that are also similar to the TRD (proof in Appendix):
Lemma 2.
lem:RL In expectation, an RL agent learning via the CL update rule follows:
@@
where $q^{\pi}_{a}$ is the action-value of $a$, and $v^{\pi}$ is the value of policy $\pi$.
With this notation, it is straightforward to combine LABEL:lem:PG and LABEL:lem:RL, provided a sensible duality exists between the identical terms. Remember how Remark 1 describes a duality between the population vector of LABEL:lem:PG and the policy of LABEL:lem:RL, $\pi$. In LABEL:lem:PG, the fitness $q^{\pi}_{a}$ is the payoff that an individual of type $a$ can expect on average when encountering a random individual from the population $\pi$. In LABEL:lem:RL, the dual of this individual of type $a$ is an action $a$ sampled from the policy $\pi$, whose expected reward is the action-value $q^{\pi}_{a}$. Similarly, the dual of the population fitness $v^{\pi}$ is the policy-value. In other words, under $R_{\text{success}}$, individuals sampled from the population $\pi$ can equivalently be regarded as action samples from an RL macro-agent, whose rewards are the individuals’ payoffs. The agent maximizes the average payoff of the group via Exact Cross Learning (we call “exact” the RL algorithm that directly applies expected updates instead of updates computed from sample estimates):
@@
Proof.
Direct consequence of LABEL:lem:PG and LABEL:lem:RL. ∎
Intuitively, this macroscopic RL process can appear as a mere by-product of individuals in the population imitating others to optimize for their own success. But let us now turn our attention to organisms that may not even have a notion of individual success to optimize: honey bees.